Use the difference quotient to find the average rate of change for this function F(X) = 5-4x. If someone could help I am completely lost.
Answer by rothauserc(4718)

The difference quotient in calculus is defined as
[ f (x + h) – f(x) ] / h
now we need to calculate f(x+h) given f(x) = 5-4x, therefore
f(x+h) = 5-4(x+h) = 5 -4x -4h
now for the difference quotient
(5 -4x -4h -5 +4x) / h = -4h/h = -4